which SI prefix multiplication of fraction designation is incorrect
1. deka 101
2. Hecto 102
3.mega 109
4. micro 10-6
Answer #1

1)

1 deca meter = 10^1 m

2)

1 hecto meter = 10^2 m

3)

1 mega meter = 10^6 m

This is incorrect

4)

1 micro meter = 10^-6 m

Answer: 3